Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Brief answer to selected problems in HW11

1. Section 5.5: problem 79: All three are correct. One can also verify using trigonometric identities cos^{2} θ = 1 − sin^{2} θ = \frac{1+cos 2θ}{2}.

2. Section 5.6: problem 112: use the change of variable u = 1 − x. problem 117: use the change of variable u = x − c.